OpenStudy (anonymous):

Please help T(I)=0.0307I Estimate l if T(I)=3070

OpenStudy (anonymous):

i got 0.00001 but it keeps saying that my answer is wrong

OpenStudy (yttrium):

That is right actually. :))

OpenStudy (anonymous):

okay, thank you @Yttrium

OpenStudy (yttrium):

No problem. :)

OpenStudy (anonymous):

\(f(I) = 0.0307 * I\) \(\frac {f(I)} {0.0307} = I\)